OpenStudy (awesome781):

A 2 kg mass moving to the right makes an elastic head-on collision with a 4 kg mass moving to the left at 4 m/s. The 2 kg mass reverses direction after the collision and moves at 3 m/s. The 4 kg mass moves 1m/s to the left. What was the initial velocity of the 2 kg mass?

OpenStudy (fifciol):

 m1v1-m2v2=-m1v1'-m2v2' 2*v1-4*4=-2*3-4*1

OpenStudy (awesome781):

I am not getting the correct answer

OpenStudy (fifciol):

I get 3m/s

OpenStudy (awesome781):

But I got 9m/s

OpenStudy (awesome781):

I figured out my mistake ....ur right.